How much water given 1M HCl and 1M Tris (pKa = 8.1) would you need to combine to make 100mL of a Tris-HCl buffer at a pH of 7.2?
Answer #1

The molar ratio of the buffer components is calculated:

n Salt / n Acid = 10 ^ (pH - pKa) = 10 ^ (7.2 - 8.1) = 0.125

It has:

i) n Salt - 0.125 * n Acid = 0

ii) n Salt + n Acid = 0.1 M * 0.1 L = 0.01 mol

System of equations is applied and you have:

n Salt = 0.0011 mol

n Acid = 0.0089 mol

The volume of Tris, HCl and water required to prepare the buffer is calculated:

V Tris = n Salt * 1000 / M = 0.0011 * 1000/1 M = 1.1 mL

V HCl = 0.0089 * 1000/1 = 8.9 mL

V H2O = 100 - 1.1 - 8.9 = 90 mL

If you liked the answer, please rate it positively, you would help me a lot, thanks.